Week of 171023

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Kate (chair, DB, WLCG), Julia (WLCG), Maarten (ALICE, WLCG), Gavin (computing), Cedric (ATLAS), Andrea M. (MW, storage), Rob (LHCb)
  • remote: Jens (NDGF), John (RAL), Andrew (NLT1), Di Qing (TRIUMF), Pepe (PIC), Kyle (OSG), Marcelo (CNAF), Xin (BNL), David B (IN2P3), Chi Hsun (ASGC), David (FNAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Normal activity + HI overlay jobs and spillover tests
      • Continuously >250k of running slots of jobs
      • Huge activity at T0 due to very good efficiency of LHC
    • Problems
      • No major issues
Xin reported that in BNL about half of ATLAS jobs have status "in transfer" instead of running. Cedric will check if this is expected

  • CMS reports ( raw view) -
    • continued high utilization of the system
    • mostly finishing HGCAL TDR samples + MC17 geant processing
    • Tier-0 under pressure on CPUs.
      • building CPU backlog since 1 week
      • buffers at high utilization, but no immediate problem
      • studying possible solutions, also in view of a similar 2018 running
      • transfers backlog stabilized if not improved
        • CNAF going to be ok soon, KIT still increasing backlog

  • ALICE -
    • Low to very high activity on average
      • Low activity due to CERN resources draining for kernel security patch (done)
    • CERN: more followup of CASTOR instabilities.
      • On Thu very few reco jobs succeeded and reco was then stopped altogether.
      • Team ticket GGUS:131217 was opened for the CASTOR team.
      • A new problem was discovered in the operation of the CASTOR Ceph gateways.
      • The situation was cleaned up and reco was then resumed.
      • We will have a meeting with the CASTOR team in the coming days.
    • CERN: EOS crashes in the night from Fri to Sat.
      • Due to some bug in the authentication module.
      • Not clear yet what triggered the bug.

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data processing and user analysis
      • pre-staging approx 50% complete, progressing well
    • Site Issues
      • T0:
        • NTR
      • T1:
        • NTR

Sites / Services round table:

  • ASGC: ntr
  • BNL: NTR
  • CNAF:
    • CMS Xroot transfers problem - under investigation
    • Transfers failing from INFN to Valencia University under investigation (it's believed to be an ipv6 issue)
  • EGI:nc
  • FNAL: ntr
  • IN2P3: the site will be in maintenance downtime on November 28th, a Tuesday. As usual details will be available one week before the event.
  • JINR: Tape robot is slightly unstable for the last few weeks. Vendor is fine tuning the system with the replacement of faulty parts.
  • KISTI: nc
  • KIT: nc
  • NDGF: ntr
  • NL-T1: ntr
  • NRC-KI: nc
  • OSG: ntr
  • PIC: CMS transfer issue between PIC and JINR is under investigation.
    • On Nov 7th half day downtime will take place to update dCache and tape library
  • RAL: NTR
  • TRIUMF: The new cluster with initial 768 cores at new data centre is in production now
    • We are using Docker container with SL7 OS, inside the container SL6 is being used.
    • The docker instances join an HTCondor pool, the CE is ARC CE.

  • CERN computing services:
    • The main CREAM CEs on LSF will be set inactive for submission tomorrow (24 October) pending retirement shortly afterwards. The EGI monitoring was disabled last week. Please send all Grid work to HTCondor CEs (except ATlasT0 backfill). OTG:0040489

  • CERN storage services:
    • FTS: FTS upgrade to v 3.7.6, tomorrow 24th Oct from 10 to 11 CEST (OTG:0040378)
    • EOSALICE: Crashed multiple times in the night of Friday to Saturday:
      • Instance suffered also from a (supposedly fixed) bug that truncates the instance's state which prevented the automatic restart to work, and required manual intervention early in the morning.
      • Identified some issues in the Authentication Plugin, being investigated.
  • CERN databases: ntr
  • GGUS: ntr
  • Monitoring:
    • Final reports for the September 2017 availability sent around
  • MW Officer: NTR
  • Networks: nc
  • Security: nc

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2017-10-23 - TommasoBoccali
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback